OpenStudy (anonymous):

Using the following equation, find the center and radius of the circle. You must show all work and calculations to receive credit. x2 −2x + y2 − 6y = 26

OpenStudy (anonymous):

[(x^2 - 2x + 1) + (y^2 - 6x + 9) = 26 + 1 + 9\] now you need to complete factor both quadratics \[(x - 1)^2 + (x - 3)^2 = 6^2\] so the general form is \[(x - h)^2 + (y - k)^2 = r^2\] so the centre is (h, k) and radius is r you need to match this to your factored equation to find the center and radius if you match \[(x -h)^2 + (y - k)^2 = r^2 ...with...(x - 1)^2 + (y - 3)^2 = 6^2\] h = 1, k = 3 so the center is (1, 3) and the radius is 6